OpenStudy (anonymous):

What is the additive inverse of the complex number 9 – 4i? A. –9 – 4i B. –9 + 4i C. 9 – 4i D. 9 + 4i

OpenStudy (vishweshshrimali5):

Additive inverse of a complex number + complex number = 0

OpenStudy (vishweshshrimali5):

Now, can you figure it out ?

OpenStudy (anonymous):

I'm guessing it would be between D and A ?

OpenStudy (vishweshshrimali5):

Try adding each of them to 9-4i. Which would result in 0 ?

OpenStudy (anonymous):

None of them will result in 0. They either come out to 5, -5 or 13.

OpenStudy (vishweshshrimali5):

No, see: (A) 9 - 4i - 9 - 4i = -8i (B) -9 + 4i + 9 - 4i = 0 (C) 9 - 4i + 9 - 4i = 18 - 8i (D) 9 + 4i + 9 - 4i = 18 So can you see which option is right ?

OpenStudy (anonymous):

Oh okay. I did the math wrong. Thanks :D

OpenStudy (vishweshshrimali5):

No problem :D

OpenStudy (anonymous):

Can you help me on one more problem?

OpenStudy (vishweshshrimali5):

Yeah sure

OpenStudy (anonymous):

I will inbox you it.

OpenStudy (vishweshshrimali5):

Ok